OpenStudy (anonymous):

When a product is raised to a power, it is equal to the product of each factor raised to that power. Show that this rule does not apply for a sum or difference to a power using the problem (5 – 3)2. You may indicate an exponent in your answer with ^. For example, 3x2 as 3x^2.

OpenStudy (phi):

the first statement **a product is raised to a power, it is equal to the product of each factor raised to that power. Show that this rule does not apply for a sum or difference to a power using the problem (5 – 3)<sup>2</sup>. You may indicate an exponent in your answer with ^. For example, 3x<sup>2</sup> as 3x^2. ** means, for example \[ (3 \cdot 5)^2 = 3^2 \cdot 5^2 \] as a check, note that 3*5= 15 \[ 15^2= 225 \] also, \[3^2 * 5^2 = 9 * 25= 225 \] it works!

OpenStudy (phi):

they want you to show **Show that this rule does not apply for a sum or difference to a power ** does not work. in other words \[ (5-3)^2 ≠ 5^2 - 3^2 \] that is not hard to do. simplify both sides, and show they are not the same number